OpenStudy (anonymous):

How do you find Arctan -sqrt3 without a calculator?

OpenStudy (anonymous):

Not sure what any of that means? This is 11th grade Trig; we've never covered any of that before?

OpenStudy (anonymous):

LOL shes referring to calculus stuff

myininaya (myininaya):

\[y=\arctan(-\sqrt{3}) \] \[\tan(y)=-\sqrt{3}=\frac{-\frac{\sqrt{3}}{2}}{\frac{1}{2}}\]